r is the midpoint of line segment pt, and q is the midpoint of line segment pr. If s is a point between r and t such that the length of segment qs is 10 and the length of segment ps is 19, what is the length of segment st?
Answer by RAY100(1637)

a sketch is very helpful
ours is just a line
point p is leftmost, point t is rightmost
r is midpoint
s is random point between r & t
Given PS=19, and qs-10
ps-qs=pq from sketch
19-10=9=pq,,,,,but pq=qr(from midpoint) therefore,,,qr=9
qs-qr=rs from sketch
10-9=1=rs,,,but rt=18,,,therefore st=18-1=17
check
p---9---q--9--r--1-s---17-----t
qs=10 ok
ps=19 ok